Marginal revenue. A defense attorney charges her client
$4000 plus $120 per hour. The formula R  120n  4000
gives her revenue in dollars for n hours of work. What is
her revenue for 100 hours of work? What is her revenue
for 101 hours of work? By how much did the one extra
hour of work increase the revenue? (The increase in
revenue is called the marginal revenue for the 101st hour.)

Answer by stanbon(75887)

A defense attorney charges her client $4000 plus $120 per hour.
The formula R = 120n + 4000 gives her revenue in dollars for n hours of work.
What is her revenue for 100 hours of work?
What is her revenue for 101 hours of work?
By how much did the one extra hour of work increase the revenue?
---------------
R(100) = 120*100 + 4000
R(100) = 1200 + 4000
R(100) = 5200
--------------
R(101) = 120*101 + 4000
R(101) = 5200 + 120 = 5320
---------------------------------
Each additional hour increases her revenue by $120.00.
---------------------------------
Notice that the 120 is the slope of the function.
margin for business is the same as slope for mathematics.
==================================
Cheers,
Stan H.